OpenStudy (anonymous):

Please help!!!!! Solve for x in terms of y. y=(10^x-10^-x)/(8)

OpenStudy (kainui):

You can rewrite \[10^{-x}=\frac{1}{10^x}\] That might help. Instead of solving for 10^x, maybe just temporarily replace \[10^x=u\] that way you're only solving for u: \[8y=u+\frac{1}{u}\] I hope that makes it easier. Think about it a little, try to figure out what I'm saying if it doesn't make sense. Try it out!

OpenStudy (anonymous):

Thank you!